# Quantum chaos in $\mathcal{PT}$-symmetric quantum systems

### v2

@@ 4.3 Parameter dependence
  • $\lambda\to 0$: This limit signifies the Hermitian limit of the problem where only two regimes have been identified with the CLSR values 0.5 and 0.57 corresponding to the integrable and chaotic regimes, respectively. These correspond to the PT-symmetric integrable regime and PT-symmetric chaotic regime for the general $(\lambda\neq 0)$ case. It can be seen from Fig. 6 that the value of $K$ for the transition between the two phases in the Hermitian limit persists even for small values of $\lambda$ for finite values of $\hbar$.
  • $\hbar\to 0$: This limit maps the quantum problem to its classical counterpart where again, there exists a phase transition from the integrable phase to a chaotic phase. As can be seen in Fig. 6, as $\hbar\to 0$, the transition from integrable-PT symmetric to chaotic-PT symmetric is happening at $K\approx 1$, which is where the classical transition occurs [56]. This observation is also in agreement with the transition observed for the Hermitian quantum kicked rotor [57] as $\hbar\to 0$. Further, we observe the shrinking of the chaotic-PT symmetric regime as $\hbar\to 0$. From the Hermitian case, we know that the chaotic-PT symmetric regime will always exist on the $\lambda=0$ line. However, we postulate that in the limit as $\hbar\to 0$, the only phase present in the $\lambda\neq 0$ space will be the chaotic-PT unbroken phase, i.e, an infinitesimal value of $\lambda$ will be sufficient to break PT symmetry, leading to chaotic nature in this limit.
Figure 6: The above figure provides a comparison between the observed PT-symmetric chaotic $\to$ PT-symmetry broken chaotic regime transition probed by the CLSR and the accompanying LE values for the same set of parameters for $K$ and $\lambda$. These have been plotted for various values of $\hbar_{\text{eff}}$ that span three orders of magnitude. We observe that this transition observed by the CLSR is accompanied by a peak in the LE. Note that different values of $K$ have only been used to ensure the transition is properly captured, since, as previously stated, the $\mathcal{PT}$-symmetric chaotic regime shrinks as $\hbar_{\text{eff}}\to 0$.
Figure 7: The above figure plots the transition from the PT-symmetric integrable regime to the PT-symmetric chaotic regime for different system sizes that increase by a factor of $2$. $\hbar=0.2$ and $K=30$ for all the plots. It can be seen that the transition, as obtained from the CLSR gets sharper as the system size ($N$) increases. There is a slight shift in the value of $\lambda$ at which the transition occurs and this shift appears to decrease with successively increasing values of $N$. The peak in the LE also appears to remain at a nearly fixed value of $\lambda$ as the system size is varied.
@@ 5 Conclusions
We have performed a comprehensive numerical study of the non-Hermitian kicked rotor demonstrating, for the first time (to our knowledge), the presence of integrability and $\mathcal{PT}$ symmetry breaking transitions in a driven quantum model. The model we have studied exhibits three phases, i) A $\mathcal{PT}$ symmetric integrable phase, ii) A $\mathcal{PT}$ symmetric chaotic phase and iii) A $\mathcal{PT}$ symmetry broken chaotic phase, thus demonstrating that $\mathcal{PT}$ symmetry breaking is a sufficient condition for the onset of chaos. We have characterized these phases and the transitions between them by calculating the complex level spacing ratio (CLSR) and the out of time ordered correlator (OTOC).
The CLSR phase diagram (see Fig. 2) elucidates several properties of the PTKR model and the intricate relation of $\mathcal{PT}$-symmetry breaking and chaos. Foremost, it shows that the CLSR is a viable diagnostic to differentiate the several phases in our system for both the non-Hermitian and Hermitian cases. One of the main observations is the absence of an integrable, $\mathcal{PT}$ symmetric phase in the phase diagram as determined from the complex level spacing ratio $\langle r\rangle$, which implies the sufficiency of $\mathcal{PT}$ symmetry breaking for the setting in of chaos. We also made a couple of other interesting observations, which, while not very clearly understood at this stage, motivated further work. The first is that a transition from between the $\mathcal{PT}$ integrable and chaotic phase can be effected by varying only the non-Hermiciticy parameter $\lambda$. This is quite intriguing since, naively, one might expect that such a transition would necessarily require varying the kicking strength, such as in the purely Hermitian system. The second observation is the appearance of a peak in the Lyapunov exponent as a function of the non-Hermiticity parameter $\lambda$ inside the $\mathcal{PT}$ symmetry broken phase close to the transition. A broad peak in the Lyapunov exponent has been observed in a Hermitian Bose-Hubbard in the vicinity of a quantum phase transition [32] but further investigation is required to determine whether or not there is any connection between the aforementioned observation and ours.
@@ A.1 Modifying the Hamiltonian
The Hamiltonian of a $\mathcal{PT}$ symmetric kicked rotor is conventionally chosen to have the following form for the potential $V(\theta)$.
$V(\theta)=K(\cos{\theta}+i\lambda\sin{\theta})$ ) ) (16)
@@ A.2 Accounting for degeneracies due to symmetries
In our case, defining the Floquet operator and retrieving the eigenvalues results in a large number of unidentified symmetries and near symmetries, causing a peak to exist near $\Delta E$ (level spacing) $=0$ in the distribution, as can be seen in figure 8. To improve the statistics and get rid of such near symmetries, we modify the kinetic energy term in the Hamiltonian. When written in the angular momentum basis, the kinetic energy is diagonal with its diagonal elements being $-\dfrac{\hbar^{2}l^{2}}{2m}$ where $l$ is the angular momentum eigenvalue and $m$ the moment of inertia. To break such symmetries, we add randomness to the masses of each diagonal term individually by redefining the diagonal elements as $-\dfrac{\hbar^{2}l^{2}}{2(m+\delta m)}$. For our calculations, $m=1$ and the $\delta m$’s are chosen randomly from a uniform distribution over the interval $[0,10^{-3}]$ for each diagonal element. This has the desired effect while still preserving the transition from the integrable to the chaotic phase. Figure 8 provides an example of this in the Hermitian case. The Hermitian version of the PTKR model has $\mathcal{P}$ and $\mathcal{PT}$ as symmetries of the system. The plot is of the Wigner-Dyson level spacing distribution.
@@ A.3 Unfolding procedure
Whether using real or complex eigenvalues, the unfolding process is crucial to getting a faithful representation of the energy level spacing distribution. The energy level spacing has the dimensions of energy. Wigner’s analysis of the level spacing distributions involved calculations over a large ensemble of $2\times 2$ matrices with particular symmetries. Suppose that this is also true for the distribution obtained from an $N\times N$ random matrix with the same symmetries over its spectrum. In this case, it must be that the level spacing in different parts of the spectrum is not related or weakly related. Consequently, in order to compare level spacing across distinct sections of the spectrum, it becomes necessary to work with dimensionless quantities. These quantities can be obtained by dividing the level spacing by the mean level spacing in a small vicinity surrounding them. Such a procedure is known as unfolding.
@@
A.4 Averaging over intervals of $K$
In order to make the level spacing ratios and distributions robust to any kind of special high symmetry $K$ values and to obtain better statistics at a much lower computational cost, we average over intervals of $K$. For kicking strengths far away from the transition points, the behavior of the energy level spacing ratio and distribution does not change much at all. Further, since the folding of the eigenvalue spectrum depends greatly on the kicking strength, there can be resonances in the quantum mechanical model reminiscent of those in the classical model. This can cause the suppression of the chaotic nature of the system. Also, rather than working with a higher dimension Hilbert space, which requires $\mathcal{O}(N^{2})$ time, we can simply combine the eigenvalues of nearby kicking strengths, which only takes $\mathcal{O}(N)$ time. Note that because the magnitude of eigenvalues can differ for different kicking strengths, we normalize the spectra for separate kicking strengths so that it lies in the interval $[0,1]$ and then concatenate them. This technique has not been employed for the calculations that yield color plots. However, this is very effective if one wishes to calculate the CLSR or RLSR in a particular neighborhood away from the transitions.
@@ Appendix B Calculation on random matrix ensembles and new CLSR
In our work, we show the agreement of the level spacing ratio obtained from our model and that from calculations over random matrices with the same symmetries as those of the model. Some of these values have been calculated previously as well [PhysRevB.106.134202]. We perform random matrix calculations for complex universality classes GinUE, GinOE, and AI [49]. Here, we highlight the method used to perform such calculations.
@@
The complex level spacing ratio can also be defined using angles, which is the other definition provided in [50]. This definition has not been adopted by us since it fails for a real eigenspectrum where $\theta$ becomes ill-defined. However, we perform random matrix calculations for $-\langle\cos{\theta}\rangle$ ) ⟩ as well as described below.
Table 2: The above table consists of data gathered on the complex level spacing ratio for some complex universality classes (GinUE, GinOE, and $\text{AI}^{\dagger}$); bracket contains must be read as (mean value, standard deviation).
Universality class $\langle r\rangle$ $-\langle\cos{\theta}\rangle$ ) ⟩
@@ Appendix C $\mathcal{PT}$-symmetry broken phases
The quasienergy of the PTKR model is complex i.e. $\epsilon=\epsilon_{r}+i\epsilon_{i}$. When the maximum value of the imaginary part (denoted by $\alpha$) exceeds a certain threshold value $\alpha_{c}$, we assume the $\mathcal{PT}$-symmetry to be broken. The yellow shaded region is the $\mathcal{PT}$ symmetry broken region. In parts of the region near the transitions, we obtain a CLSR close to 0.739. This is the value of the CLSR one obtains for the GinOE universality class, to which the Hamiltonian belongs. This is due to the fact that the Hamiltonian is not bound by Hermiticity, and it commutes with an anti-unitary operator (namely $\mathcal{PT}$). In the table below, we show calculations of the CLSR on random matrices in the GinOE class. Additionally, we also perform calculations on matrices within GinOE that commute with $\mathcal{PT}$. These are in agreement with the value of the CLSR we get in our phase diagram.
Table 3: The above table consists of data gathered on the CLSR for the $\mathcal{PT}$ symmetry broken chaotic regime and its comparison with CLSR in the complex universality classes GinOE and its subset of matrices that commute with $\mathcal{PT}$, bracket contents must be read as (mean value, standard deviation).

### v3

@@ 4.3 Parameter dependence
  • $\lambda\to 0$: This limit signifies the Hermitian limit of the problem where only two regimes have been identified with the CLSR values 0.5 and 0.57 corresponding to the integrable and chaotic regimes, respectively. These correspond to the $\mathcal{PT}$-symmetric integrable regime and $\mathcal{PT}$-symmetric chaotic regime for the general $(\lambda\neq 0)$ case. It can be seen from Fig. 6 that the value of $K$ for the transition between the two phases in the Hermitian limit persists even for small values of $\lambda$ for finite values of $\hbar_{\text{eff}}$.
  • $\hbar_{\text{eff}}\to 0$: This limit maps the quantum problem to its classical counterpart where again, there exists a phase transition from the integrable phase to a chaotic phase. As can be seen in Fig. 6, as $\hbar_{\text{eff}}\to 0$, the transition from $\mathcal{PT}$-symmetric intergrable to $\mathcal{PT}$-symmetric chaotic is happening at $K\approx 1$, which is where the classical transition occurs [60]. This observation is also in agreement with the transition observed for the Hermitian quantum kicked rotor [61] as $\hbar_{\text{eff}}\to 0$. Further, we observe the shrinking of the $\mathcal{PT}$-symmetric chaotic regime as $\hbar_{\text{eff}}\to 0$. From the Hermitian case, we know that the $\mathcal{PT}$-symmetric chaotic regime will always exist on the $\lambda=0$ line. However, we postulate that in the limit as $\hbar_{\text{eff}}\to 0$, the only phase present in the $\lambda\neq 0$ space will be the $\mathcal{PT}$-symmtric chaotic unbroken phase, i.e, an infinitesimal value of $\lambda$ will be sufficient to break $\mathcal{PT}$-symmetry, leading to chaotic nature in this limit.
Figure 6: The above figure shows a section of the CLSR contour map for different values of $\hbar_{\text{eff}}$. It can be observed that the transition $T^{*}$ moves closer to the integrable $\to$ chaotic transition seen in the classical model for which $K\sim 1$ as $\hbar_{\text{eff}}\to 0$. Furthermore, the $\mathcal{PT}$-symmetric chaotic regime is observed to shrink in size with a decrease in $\hbar_{\text{eff}}$
@@ 5 Conclusions
We have performed a comprehensive numerical study of the non-Hermitian kicked rotor demonstrating, for the first time (to our knowledge), the presence of integrability and $\mathcal{PT}$-symmetry breaking transitions in a driven quantum model. The model we have studied exhibits three phases, i) A $\mathcal{PT}$-symmetric integrable phase, ii) A $\mathcal{PT}$-symmetric chaotic phase and iii) A $\mathcal{PT}$-symmetry broken chaotic phase, thus demonstrating that $\mathcal{PT}$-symmetry breaking is a sufficient condition for the onset of chaos. We have characterized these phases and the transitions between them by calculating the complex level spacing ratio (CLSR) and the out of time ordered correlator (OTOC).
The CLSR phase diagram (see Fig. 2) elucidates several properties of the PTKR model and the intricate relation of $\mathcal{PT}$-symmetry breaking and chaos. Foremost, it shows that the CLSR is a viable diagnostic to differentiate the several phases in our system for both the non-Hermitian and Hermitian cases. One of the main observations is the absence of an integrable, $\mathcal{PT}$-symmetric phase in the phase diagram as determined from the complex level spacing ratio $\langle r\rangle$, which implies the sufficiency of $\mathcal{PT}$-symmetry breaking for the setting in of chaos. Equivalently, this indicates the absence of a $\mathcal{PT}$-symmetry broken, integrable phase in our model. Note however that the model is integrable with broken $\mathcal{PT}$ symmetry when $\lambda=\infty$ and $m=\infty$. In this limit, the Floquet eigenvalues are simple $iK\sin\theta$ for all $\theta\in[0,2\pi)$. Thus, the eigenvalues are pure imaginary, occur in complex conjugate pairs showing that the system breaks $\mathcal{PT}$ symmetry. Moreover, since they can also be obtained exactly, the system is clearly integrable. However, it appears from our numerical study that this special $\mathcal{PT}$ symmetry broken integrable point does not extend into a phase for finite $m$ and $\lambda$. The question of whether the absence of such a phase is special to our model or more generic will require further investigation. We have also made a few other other interesting observations, which, while not very clearly understood at this stage, motivated further work. The first is that a transition from between the $\mathcal{PT}$-symmetric integrable and chaotic phase can be effected by varying only the non-Hermiciticy parameter $\lambda$. This is quite intriguing since, naively, one might expect that such a transition would necessarily require varying the kicking strength, such as in the purely Hermitian system. The second observation is the appearance of a peak in the Lyapunov exponent as a function of the non-Hermiticity parameter $\lambda$ inside the $\mathcal{PT}$-symmetry broken phase close to the transition. A broad peak in the Lyapunov exponent has been observed in a Hermitian Bose-Hubbard in the vicinity of a quantum phase transition [32] but further investigation is required to determine whether or not there is any connection between the aforementioned observation and ours.
@@ A.1 Modifying the Hamiltonian
The Hamiltonian of a $\mathcal{PT}$-symmetric kicked rotor is conventionally chosen to have the following form for the potential $V(\theta)$.
$))\displaystyle V(\theta)=K(\cos{\theta}+i\lambda\sin{\theta})$ (10)
@@ A.2 Accounting for degeneracies due to symmetries
In our case, defining the Floquet operator and retrieving the eigenvalues results in a large number of unidentified symmetries and near symmetries, causing a peak to exist near $\Delta E$ (level spacing) =0=0= 0 in the distribution, as can be seen in figure 7. To improve the statistics and get rid of such near symmetries, we modify the kinetic energy term in the Hamiltonian. When written in the angular momentum basis, the kinetic energy is diagonal with its diagonal elements being $-\dfrac{\hbar^{2}l^{2}}{2m}$ where $l$ is the angular momentum eigenvalue and $m$ the moment of inertia. To break such symmetries, we add randomness to the masses of each diagonal term individually by redefining the diagonal elements as $-\dfrac{\hbar^{2}l^{2}}{2(m+\delta m)}$. For our calculations, $m=1$ and the $\delta m$’s are chosen randomly from a uniform distribution over the interval $[0,10^{-3}]$ for each diagonal element. This has the desired effect while still preserving the transition from the integrable to the chaotic phase. Figure 7 provides an example of this in the Hermitian case. The Hermitian version of the PTKR model has $\mathcal{P}$ and $\mathcal{PT}$ as symmetries of the system. The plot is of the Wigner-Dyson level spacing distribution.
@@ A.3 Unfolding procedure
Whether using real or complex eigenvalues, the unfolding process is crucial to getting a faithful representation of the energy level spacing distribution. The energy level spacing has the dimensions of energy. Wigner’s analysis of the level spacing distributions involved calculations over a large ensemble of $\times 2$ 2 matrices with particular symmetries. Suppose that this is also true for the distribution obtained from an $N\times N$ random matrix with the same symmetries over its spectrum. In this case, it must be that the level spacing in different parts of the spectrum is not related or weakly related. Consequently, in order to compare level spacing across distinct sections of the spectrum, it becomes necessary to work with dimensionless quantities. These quantities can be obtained by dividing the level spacing by the mean level spacing in a small vicinity surrounding them. Such a procedure is known as unfolding.
@@ Appendix B Calculation on random matrix ensembles and new CLSR
In our work, we show the agreement of the level spacing ratio obtained from our model and that from calculations over random matrices with the same symmetries as those of the model. Some of these values have been calculated previously as well [55]. We perform random matrix calculations for complex universality classes GinUE, GinOE, and AI [53]. Here, we highlight the method used to perform such calculations.
Figure 8: The above figure represents depicts the spread of the quasienergies in the three different regimes. Top left : $\mathcal{PT}$ symmetric integrable, Bottom left : $\mathcal{PT}$ symmetric chaotic, Right : $\mathcal{PT}$ symmetry broken chaotic.
@@
The complex level spacing ratio can also be defined using angles, which is the other definition provided in [54]. This definition has not been adopted by us since it fails for a real eigenspectrum where $\theta$ becomes ill-defined. However, we perform random matrix calculations for $)-\langle\cos{\theta}\rangle$ as well as described below.
Table 2: The table below consists of data gathered on the complex level spacing ratio for some complex universality classes (GinUE, GinOE, and $\text{AI}^{\dagger}$); bracket contains must be read as (mean value, standard deviation).
Universality class $\langle r\rangle$ $)-\langle\cos{\theta}\rangle$
@@ Appendix C $\mathcal{PT}$-symmetry broken phases
The quasienergy of the PTKR model is complex i.e. $\epsilon=\epsilon_{r}+i\epsilon_{i}$. When the maximum value of the imaginary part (denoted by $\alpha$) exceeds a certain threshold value $\alpha_{c}$, we assume the $\mathcal{PT}$-symmetry to be broken. The yellow shaded region in fig.9 is the $\mathcal{PT}$ symmetry broken region. In parts of the region near the transitions, we obtain a CLSR close to 0.739. This is the value of the CLSR one obtains for the GinOE universality class, to which the Hamiltonian belongs. This is due to the fact that the Hamiltonian is not bound by Hermiticity, and it commutes with an anti-unitary operator (namely $\mathcal{PT}$). In the table 3, we show calculations of the CLSR on random matrices in the GinOE class. Additionally, we also perform calculations on matrices within GinOE that commute with $\mathcal{PT}$. These are in agreement with the value of the CLSR we get in our phase diagram.
Figure 9: Left: The maximum of imaginary part $\alpha$ of the eigenspectrum for varying values of the kicking strength $K$ and the non-hermiticity $\lambda$, computed for $N=4096$. The dotted lines $T_{1}$, $T_{2}$, and $T_{3}$ represent the transition across the $\mathcal{PT}$-symmetric chaotic phase to the $\mathcal{PT}$ symmetry broken chaotic phase, $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetry broken chaotic phase, and the $\mathcal{PT}$-symmetric integrable phase to the $\mathcal{PT}$-symmetry broken chaotic phase. Right: The color plot highlights the region of the phase diagram for which the CLSR is close to 0.5 which corresponds to the $\mathcal{PT}$ symmetric integrable phase.
Table 3: The table below consists of data gathered on the CLSR for the $\mathcal{PT}$ symmetry broken chaotic regime and its comparison with CLSR in the complex universality classes GinOE and its subset of matrices that commute with $\mathcal{PT}$, bracket contents must be read as (mean value, standard deviation).
@@
Appendix D Unnormalized OTOC and Norm Growth
Figure 10: The OTOC $C(t)$ and state norm $\mathcal{N}(t)$ vs $t$ across the transition along Left $T_{2}$ and Right $T_{1}$ shown in Fig. 2, calculated for system size $N=2^{14}$. For $T_{1}$ (Left), the values $(\lambda,K)=\{(0.001,0.1),(0.002,0.456),(0.003,1.135)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken integrable phase, while the values $(\lambda,K)=\{(0.006,5.734),(0.01,15.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. For $T_{2}$ (right), the values $(\lambda,K)=\{(10^{-5},30),(10^{-4},30),(10^{-3},30)\}$ correspond to the $\mathcal{PT}$-symmetry unbroken chaotic phase, while the values $(\lambda,K)=\{(0.005,30),(0.001,30.0)\}$ correspond to the $\mathcal{PT}$-symmetry broken chaotic phase. In both cases, the $\mathcal{PT}$-symmetric broken phase displays a late-time exponential growth of the OTOC and the state norm.
In the $\mathcal{PT}$-symmetry broken phase, the eigenspectrum consists of complex eigenvalues. As a result, the time-evolved state $|\Psi(t)\rangle=U(t)|\Psi\rangle$ no longer remains normalized (see Fig. 10). The late-time growth of both the OTOC and the state norm depends on the eigenspectrum — in particular, it depends strongly on the maximum of the imaginary part of the eigenvalues of the Hamiltonian. Let us denote this maximum value by $\alpha$. At late times, the evolution is dominated by the eigenvalue with the largest imaginary part:
Consequently, the late-time norm $\mathcal{N}(t)=\langle\Psi(t)|\Psi(t)\rangle$ scales as:
$\mathcal{N}(t\rightarrow\infty)\propto e^{2\alpha t}\,.$ (14)
To analyze the behavior of the OTOC, we expand the squared commutator $-[\hat{p}(t),\allowbreak\hat{p}(0)]^{2}$ as
[…]
Hence, the unnormalized OTOC exhibits exponential growth:
$C(t\rightarrow\infty)\propto e^{4\alpha t}\,$ (17)
This motivates the definition of a normalized OTOC:
$\tilde{C}(t)=e^{-4\alpha t}C(t)\,.$ (18)
which removes the exponential divergence associated with the non-unitary dynamics. The normalized OTOC $\tilde{C}(t)$ thus captures only the growth associated with the chaotic nature of the system.
Since our calculations are performed at finite times, subleading eigenvalues can still contribute to the dynamics. Therefore, the normalization factor $\alpha$ is determined by fitting the late-time growth of $C(t)$ to an exponential form.